# Paying Attention to Deflections: Mining Pragmatic Nuances for Whataboutism Detection in Online Discourse

### v1

@@ 4.3 Mining negatives with attention
Given the centrality of topical redirection in whataboutism, any representation distinguishing it requires an understanding of the pragmatic context, namely the antecedent QUD. We thus model each comment by training on tuples comprising the comment (which we seek to classify), along with comments each from the w and nw classes. For our collection of YouTube comments (TQ), 2 comments are selected from the same video as ; for TQ, from the same thread as the original tweet. This hyperparameter controls context incorporation during training. We reason that such tuples can better capture semantic redirection within the ambient context of the social topic, than embeddings based solely on global distributional semantics lenci2022. With this setup121212The language model , which provides the representation of and contextual comments, can be viewed as a discrete parameter of the complete approach, where one model may be swapped out for another., the most dissimilar examples can be identified using a context tuple. Then, combining them with a representation of will encode pragmatic shifts through an implicit grounding of the text within the surrounding social discourse.131313In effect, we surmise that the pragmatic act is grounded not just in the context of immediately surrounding texts, but in a sampling of the entire social discourse about that topic. In spirit, this is similar to the idea of visually grounding objects through cross-modal attention ilinykh2022attention.
Specifically, we apply a Transformer encoder with layers and attention heads vaswani2017attention to incorporate cross-attention between the elements of the context tuple. We then extract the cross-attention map from the encoder’s final layer for a similarity matrix. After identifying the example with the highest numeric value, we concatenate its text with . This combined embedding is fed into a multi-layer perceptron for final classification. Termed Mining Negatives with Aattention (MINA), this method employs cross-attention scores to mine the most pragmatically dissimilar examples in the discourse surrounding . Fig. 3 is an in situ illustration of its complete architecture. Ablation studies (§ 5) dictate our configuration of 2 encoder layers, 32 attention heads, SBERT embeddings as input, and a context size of 1.
@@ 5 Results and analysis
Baseline models.
Table 2 contains the results of the baseline transfer-learning models, models based on semantic similarity, and MINA. While the baseline models perform reasonably well on Twitter, their performance on TQ is underwhelming: BERT achieves higher precision but lower recall, while RoBERTa shows the opposite behavior. It is thus evident that models adept at propaganda detection have room for improvement in identifying whataboutism, especially given their tendency to assign the majority label nw to test instances. This issue is particularly relevant for real-world data, where – much like our collections – nw is likely to dominate as the majority class.
Semantic similarity models.
@@
MINA.
Even while employing contextual embeddings, the inability of the above models to effectively detect whataboutism suggests a limitation in capturing the kind of pragmatic nuances illustrated earlier (Fig. 1, 2). MINA, however, leverages cross-attention scores across comments within a topic to model pragmatic context. To demonstrate the efficacy of our approach, we assess the worst-performing models from previous categories, focusing on their F1 scores for the target class. On TQ, of RoBERTa increases by 22%, rising from 0.3 to 0.52, while SBERT shows an even more remarkable improvement of 51%, from 0.07 to 0.58: a 10% boost over the next best result.141414For comparison, the SemEval-2020 task on propaganda detection dasanmartino2020semeval saw the best performance in detecting whataboutism, straw man arguments, and red herring fallacies achieve = 0.269 (the task grouped these three types of fallacies due to insufficient data; there were only 76 instances of whataboutism, for instance). When evaluated on TQ, SBERT with MINA outperforms other models, showing a 6% improvement over vanilla SBERT and a 4% improvement over RoBERTa and SimCSE. RoBERTa initially exhibits high recall but low precision due to being predisposed to irony. However, MINA often flips the output when irony is present in comments but not in other context tuple elements.
The significant improvements we observe with MINA are reflected in its effects on target class separability. Comparing the t-SNE hinton2002stochastic visualisations in Fig. 3(a) and Fig. 3(b), it becomes clear that while fine-tuning on a moderate amount of task-specific training data has limited utility, contrasting with MINA makes whataboutism instances easier to discriminate.
Ablation experiments.
@@
(a) Varying , with (): |
(attention heads) |
32 |
0.52 |
0.71 |
64 |
0.52 |
0.70 |
128 |
0.51 |
0.70 |
384 |
0.48 |
0.69 |
(b) Varying , with (): |
(context size) |
1 |
0.51 |
0.71 |
2 |
0.50 |
0.70 |
3 |
0.51 |
0.67 |
4 |
0.49 |
0.69 |
(c) Varying , with (): |
(language model) |
sbert |
0.52 |
0.71 |
roberta Irony |
0.51 |
0.70 |
A comparison of mining strategies
@@
We compare MINA’s negative sample mining with two common strategies in recent contrastive mining literature: cosine similarity-based mining (e.g., wang2021dense) and random sampling jiang2021improving; xu2022negative. Table LABEL:tab:mining-comparison shows MINA’s superior result, while the use of cosine measure is akin to random sampling for tuple construction. This further supports our conjecture that conventional measures of semantic similarity or contrast are inadequate for capturing pragmatic differences.
@@
Everything is included. Due to the dynamic nature of social media platforms and unforeseeable changes to a platform’s policies, however, there is no guarantee that future use of the data collection script will be able to collect the same data again, even if the comments/tweets continue to remain on YouTube/Twitter.
@@ A.4 Legal & ethical considerations
No. However, the dataset does not contain (and indeed, even the data collection process does not include) any names or user handles.
@@
TQ+ does not contain any personal data of the authors of the collected comments. Further, it does not contain information that might be considered sensitive or confidential.
@@ Appendix B Proof of near-orthogonality
Formal proof of the following statement (§ 4.2):
“Cosine similarity may present challenges in high-dimensional vector spaces, as it becomes highly probable that any two vectors are nearly orthogonal.”

### v2

@@ 4.3 Mining negatives with attention
Given the centrality of topical redirection in whataboutism, any representation distinguishing it requires an understanding of the pragmatic context, namely the antecedent QUD. We thus model each comment by training on tuples comprising the comment (which we seek to classify), along with comments each from the w and nw classes. For our collection of YouTube comments (TQ), the 2 comments are selected from the same video as ; and for TQ, from the same thread as the original tweet. This hyperparameter controls context incorporation during training. We reason that such tuples can better capture semantic redirection within the ambient context of the social topic, than embeddings based solely on global distributional semantics Lenci et al. (2022). With this setup131313The language model , which provides the representation of and contextual comments, can be viewed as a discrete parameter of the complete approach, where one model may be swapped out for another., the most dissimilar examples can be identified using a context tuple. Then, combining them with a representation of will encode pragmatic shifts through an implicit grounding of the text in the surrounding social discourse.141414In effect, we surmise that the pragmatic act is grounded not just in the context of immediately surrounding texts, but in a sampling of the entire social discourse about that topic. This is similar in spirit to visually grounding objects through cross-modal attention Ilinykh and Dobnik (2022).
Specifically, we apply a Transformer encoder with layers and attention heads Vaswani et al. (2017) to incorporate cross-attention between the elements of the context tuple. We then extract the cross-attention map from the encoder’s final layer for a similarity matrix. After identifying the example with the highest numeric value, we concatenate its text with . This combined embedding is fed into a multi-layer perceptron for final classification. Termed mining negatives with attention (mina), this method employs cross-attention scores to mine the most pragmatically dissimilar examples in the discourse surrounding . Fig. 3 is an in situ illustration of its complete architecture. Ablation studies (§5) dictate our chosen configuration of 2 encoder layers, 32 attention heads, SBERT embeddings as input, and a context size of 1.
@@ 5 Results and analysis
Table 2 displays the results of baseline transfer-learning on pretrained Transformer-based models, models based on semantic similarity, and mina.
Baseline models:
The baseline models perform reasonably well on Twitter, but not on TQ: BERT achieves higher precision but lower recall, while RoBERTa shows the opposite behavior. It is thus evident that models adept at propaganda detection have room for improvement in identifying whataboutism, especially given their tendency to assign the majority label nw to test instances.
Semantic similarity models:
@@
Mining negatives with attention (mina):
Even while employing contextual embeddings, the inability of the above models to effectively detect whataboutism suggests a limitation in capturing the kind of pragmatic nuances illustrated earlier (Fig. 1, 2). mina, however, is designed to leverage cross-attention scores across comments within a topic to model pragmatic context. To demonstrate its efficacy, we assess the worst-performing models from previous categories, focusing on for the target class. On TQ, of RoBERTa increases by 22%, rising from 0.3 to 0.52, while SBERT shows an exceptional improvement of 51%, from 0.07 to 0.58: a 10% boost over the next best result.151515To compare, the SemEval-2020 task on propaganda detection Da San Martino et al. (2020) saw the best performance in detecting whataboutism, straw man arguments, and red herrings, achieve = 0.269. (these three closely-related fallacies were merged into a single group due to insufficient data. When evaluated on TQ, SBERT with mina outperforms other models, showing a 6% improvement over vanilla SBERT and a 4% improvement over RoBERTa and SimCSE. RoBERTa initially exhibits high recall but low precision due to being predisposed to irony. However, mina often flips the output when irony is present in comments but not in other context tuple elements. The significant improvements we observe with mina are reflected in its effects on target class separability. Comparing the t-SNE Hinton and Roweis (2002) visualizations in Fig. 4(a) and Fig. 4(b), it becomes clear that while fine-tuning on a moderate amount of task-specific training data has limited utility, contrasting with mina improves the ability to distinguish whataboutism from everything else.
(a) Varying , with (): | ||
(attention heads) | (YouTube) | (Twitter) |
32 | 0.52 | 0.71 |
64 | 0.52 | 0.70 |
128 | 0.51 | 0.70 |
384 | 0.48 | 0.69 |
(b) Varying , with (): | ||
(context size) | (YouTube) | (Twitter) |
1 | 0.51 | 0.71 |
2 | 0.50 | 0.70 |
3 | 0.51 | 0.67 |
4 | 0.49 | 0.69 |
(c) Varying , with (): | ||
(language model) | (YouTube) | (Twitter) |
sbert | 0.52 | 0.71 |
roberta Irony | 0.51 | 0.70 |
Ablation experiments:
@@
Method | (YouTube) | (Twitter) |
MINA | 0.52 | 0.71 |
Random | 0.50 | 0.67 |
Cosine-Sim | 0.50 | 0.67 |
A comparison of mining strategies:
@@
We compare mina’s negative sample mining with two common strategies in recent contrastive mining literature: cosine similarity-based mining (e.g., Wang et al., 2021) and random sampling Jiang et al. (2021); Xu et al. (2022). Table 4 shows mina’s superior result, while the use of cosine measure is akin to random sampling for tuple construction. This further supports our conjecture that conventional measures of semantic similarity or contrast are inadequate for capturing pragmatic differences.
Challenges, negative results, and insights:
In exploring several models and techniques to detect whataboutism, the path to mina, which uses negative sampling inspired by cross-attention, was marked by several unexpected outcomes. First, we experimented with the advanced pre-trained encoder DeBERTa He et al. (2021) on the collection of YouTube comments, but its performance was unexpectedly inferior to BERT, achieving an score of 0.455 compared to BERT’s 0.48.
Although modern large language models (LLMs) like Llama-2-7b, Llama-2-13b Touvron et al. (2023), and GPT-3.5-turbo Brown et al. (2020) have shown impressive capabilities in various tasks, they did not yield satisfactory results in detecting whataboutism. The limitations we observe align with those noted by Ruiz-Dolz and Lawrence (2023) in detecting fallacious argumentation. On Twitter comments, for instance, the highest score was 0.65, 4% lower than the BERT baseline (Table 2). Even with sophisticated contextual prompting, no significant improvements were observed. GPT-3.5-turbo excelled with canonical whataboutism that closely mirrored syntactic patterns found in popular media, but struggled with nuanced, rephrased instances. Consequently, these LLMs were excluded from subsequent stages of experimentation in this work.
Our experiments included incorporating video transcripts as contextual information, for which we leveraged Longformer Beltagy et al. (2020). Contrary to initial expectations, these yielded lower precision and recall across all models. Analyzing the errors revealed that comments from other users are a better predictor than the contextual information derived from the video. This observation, plus our thematic analysis (§ 3.1) and the success of mina strongly indicate that while the video remains important in setting the topic, detecting whataboutism in a comment often depends more on the broader social discourse surrounding the topic.
Lastly, we decided against batch-contrastive learning due to two potential concerns raised by Khosla et al. (2020) and Qu et al. (2021): (i) large batch sizes are vital for it to be effective, and (ii) it typically involves augmentation, which introduces noise that hinders pragmatic understanding. Further, it would still use cosine similarity on its own, which we have questioned in § 4.2.
6 Related Work
Propaganda detection:
Early propaganda detection Barrón-Cedeño et al. (2019) categorized sources based on external inputs, but concerns were raised due to noisy data and questions about epistemic bias Uscinski and Butler (2013); Marietta et al. (2015). Traditionally, whataboutism was viewed as the tu quoque fallacy Fischer (2021). Another body of work, however, views it in the broader pragmatic and dialectic framework Aikin (2008). While this has gained prominence in theoretical studies Bowell (2023), computational approaches remain confined to viewing whataboutism purely as propaganda Da San Martino et al. (2019b); Sahai et al. (2021); Baleato Rodríguez et al. (2023). Our work connects to the pragma-dialectic view of whataboutism: overlapping with propaganda but not subsumed by it.
NLP tasks have traditionally favored utilizing pretrained models like BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), fine-tuning them for specific tasks. Da San Martino et al. (2019b) accordingly approached propaganda detection, creating a corpus of news articles later used in shared tasks Da San Martino et al. (2019a, 2020), where these pretrained models achieved top results Yoosuf and Yang (2019); Morio et al. (2020). While related, our research differs substantially. They target propaganda detection, treating whataboutism only as propagandist language while the limited dataset size precludes deeper analysis. Da San Martino et al. (2020) nevertheless note the difficulty in identifying whataboutism.
Since identifying the precise textual span of propaganda is laborious and error-prone, we look to Sahai et al. (2021), who label informal fallacies in social media comments. However, they continue to treat whataboutism purely as propaganda, while also turning to transfer learning through fine-tuning the models (i.e., the approach we use as baselines).
Framing and persuasion:
Unlike earlier studies, Piskorski et al. (2023) view whataboutism as strategic framing and persuasion rather than outright propaganda, thus offering a taxonomy that aligns more closely with our study. However, instances of whataboutism are sparse in their multilingual corpus (only 25 in English), limiting its utility in our work. Additionally, fine-tuned Transformer models or ensembles, as utilized by top-performing participants, yielded remarkably low scores for whataboutism and other related categories (see Wu et al., 2023). The results of this shared task are noteworthy, yet they are confounded by ambiguity Purificato and Navigli (2023); Liu et al. (2023) and low inter-annotator agreement. Addressing these challenges is crucial for an accurate interpretation or analysis of the findings.
Modeling (dis) similarity.
Whataboutism, a discursive maneuver for diversion and centering, relates to semantic textual similarity (STS) Cer et al. (2017). Modeled as a graph, STS becomes a node classification task using graph neural networks (GNN). Huang et al. (2021), however, show that minor modifications to classical graph-based learning algorithms (e.g., Zhou et al., 2003) outperform large GNNs despite far fewer parameters and less training time. Their approach (C&S) does not require careful early stopping criteria or a large validation set, making it another competitive baseline.
STS models do not capture whataboutism’s pragmatic complexity. We address this by sharing a core intuition with contrastive learning: bring similar instances closer while pushing apart dissimilar ones Weinberger and Saul (2009). Augmentation methods used in contrastive learning, however, are ill-suited for pragmatics. Instead, we opt for SimCSE Gao et al. (2021). While hard negative mining has proven effective Xiong et al. (2021), the use of attention weights to model pragmatic differences is a novel contribution of this work. Here, notable work includes Le et al. (2018), who integrate attention weights with syntax, and Yamagiwa et al. (2022), who fuse self-attention matrices with word-mover’s distance to obtain promising STS results.
7 Conclusion
Our study underscores areas where state-of-the-art propaganda detection models often fail to distinguish propagandist use of whataboutism from valid argumentation tools or figurative language use like irony, revealing how their suitability for this task may be improved. Addressing this within the pragma-dialectic framework, we illustrate the need to develop models capable of understanding pragmatic variations. We find traditional similarity measures ineffective for this purpose. Cross-attention proves valuable instead. We thus propose a novel methodology, mina, for mining negative samples based on cross-attention, which achieves superior results in whataboutism detection and suggests its utility in other tasks that require grounding in ambient social discourse. This study is based on our contribution of two annotated datasets that can facilitate further research in related areas.
8 Limitations
Emotive topics may influence the labels assigned to specific instances. We report Fleiss’ kappa, adjusting for the possibility of chance agreement, but our approach does not take into account individual viewpoints. We thus advocate imbibing this into perspectivist research Cabitza et al. (2023).
We have shared model weights, code, and datasets for replicability. mina’s reliance on randomized tuple batches, however, may cause slight changes in retrained models in spite of extremely low variance (reported in Table 3).
The insights regarding the underwhelming performance of modern LLMs are crucial for understanding their limitations in handling tasks like whataboutism detection, where pragmatic nuances are important. In this regard, a deeper exploration of in-context learning is warranted.
Finally, we note that the TQ collections are imbalanced, with a minority target class. This, however, mirrors real-world scenarios and may not be a disadvantage after all.
Appendix A Datasheet
This datasheet is included based on the recommendations of Gebru et al. (2021). It serves to document the creation, composition, intended uses, and maintenance of the TQ+ dataset (“tu quoque and beyond”) released with this work. We hope this will facilitate its better usage, and further encourage transparency, accountability, and reproducibility.111For the sake of brevity, we have not included information that is included in the main body of this paper.
Why was the dataset created?
TQ+ was created to enable research on identifying expressions of whataboutism in social media posts made by active participating users: given an English (en-US) comment made in the specific context, identify whether the comment expresses whataboutism. Intentionally created for this task, the dataset comprises of several hot-button sociopolitical topics, where whataboutism is more likely to be employed. While there exist a few datasets for propaganda detection Da San Martino et al. (2019a), there is no corpus exclusively devoted to the study of whataboutism, which has unique sociopolitical and linguistic properties (ranging from semantics to pragmatics) that are distinct from other propagandist maneuvers. Furthermore, many who employ whataboutism are neither the creators nor the intentional disseminators of propaganda. Instead, many instances are attempts made by users to challenge or re-center a narrative (better captured in the taxonomic changes seen in Piskorski et al. (2023)). These factors distinguish TQ+ from earlier datasets.
What other tasks could the dataset be used for?
TQ+ can be used for various types of modeling or comprehension of whataboutism in social media commentary. For instance, one may study syntactic patterns and their correlation with expressions of whataboutism, or how other users react to comments that express whataboutism. It can also be used to learn from adversarial settings, wherein this dataset can be used to automatically generate whataboutism on politically divisive topics. Furthermore, we expect TQ+ to be useful for discourse analyses and other studies conducted by social scientists and media communication researchers.
Has the dataset been used for any tasks already?
This is a novel dataset. As of June 2023, it has not been used for any other task or publication.
A.1 Dataset composition
TQ+ comprises two sub-datasets: TQ, which consists of comments made by active users on YouTube videos, and TQ, which consists of Twitter posts.
Fig. 5(a) shows the distribution of these comments and their labels across the topics in the YouTube corpus.The Twitter corpus has 6 topics that are in TQ, and additionally, Tiktok ban and Trump indictment. The data distribution across all these topics, along with the distribution of the binary ground-truth labels, is shown in Fig. 5(b).
What are the instances?
For both collections, three annotators worked independently to generate the labels. Since multiple annotator judgments are obtained before the majority vote, we take cognizance of the recommendations made by Prabhakaran et al. (2021), and provide the individual annotator labels in the released datasets for flexible future use.
Are relationships between instances made explicit in the data?
@@ A.1 Dataset composition
Everything is included. Due to the dynamic nature of social media and potential policy changes, however, future use of the data collection script may not gather the same data even if the comments remain on the platform.
@@ A.4 Legal & ethical considerations
No. Neither the dataset nor the data collection process includes any names or user handles.
@@
TQ+ does not contain any personal data of the authors of the collected comments. Further, it does not contain sensitive or confidential information.
@@ Appendix B Proof of near-orthogonality
Formal proof of the statement (§ 4.2): Cosine similarity may present challenges in high-dimensional vector spaces, as it becomes highly probable that any two vectors are nearly orthogonal.
@@
Appendix C Training and Validation Loss Analysis
This appendix presents the learning curves for the models enhanced with mina, illustrating the training and validation loss over all epochs. These plots, shown in Fig. 6, depict their performances and generalization capabilities, addressing potential concerns about overfitting due to the moderate size of the TQ collections.